# scenarioparampadding=8pt, margin-top=3pt, margin-bottom=3pt, border-style=none, background-color=quotebackground

### v1

@@ 5.2. Data Collection
Interviews with Patients. We conducted interviews with each patient to delve deeply into their experiences and learn how they used MindfulDiary on a daily basis. At the end of the second and fourth weeks, we conducted 15-minute debriefing interviews with participants. Considering the characteristics of depression patients, who may struggle to focus for long periods of time, the interview session was divided into two shorter sessions.
@@
Interviews with Psychiatrists. We interviewed psychiatrists who treated the patient participants to understand how they might use the system’s data in actual clinical settings. We further gathered feedback from the psychiatrists on the opportunities and limitations of MindfulDiary, as well as suggestions for improvements.
@@ 5.3. Analysis
To explore participants’ usage patterns with MindfulDiary, we first conducted a descriptive statistics analysis. To determine any shifts in participants’ adherence over time, we examined weekly writing frequencies using a one-way repeated measures ANOVA (RM-ANOVA) with Greenhouse-Geisser correction. To gain a deeper qualitative understanding of the messages produced by MindfulDiary and interviews with patients and psychiatrists, we used open coding paired with thematic analysis (Braun and Clarke, 2006). All interviews were audio-recorded and later transcribed. The first author open-coded the interview transcripts along with the interaction log data through multiple rounds of iteration. The research team then identified patterns through discussion to formulate overarching themes.
@@ 5.4. Ethical Considerations
Our study design obtained the required approval from the Institutional Review Board (IRB) of the affiliated hospital.
Conducting this study, we are fully aware of the inherent risks associated with our research, particularly given the characteristics of participants diagnosed with MDD. To mitigate the risks, we first carefully screened participants, relying on evaluations conducted by psychiatrists. Individuals displaying heightened impulsive tendencies or harboring specific intentions towards self-harm or suicide were excluded from the study. In addition, participants were asked to take the PHQ-9 before interacting with MindfulDiary, along with an additional set of questions probing their recent attempts at self-harm or suicide. If a participant’s response to question number 9 of the PHQ-9, regarding suicidal/self-harm thoughts, scored ‘moderate or higher’ or if any recent suicide attempt was verified, the system pivoted to provide content geared towards alleviating anxiety and reducing stress rather than proceeding with the standard program. In such a case, a real-time alert was also sent to psychiatrists. Lastly, if sensitive themes frequently surfaced in a participant’s input during the study, their interactions with the program were temporarily halted. Psychiatrists subsequently re-evaluated such participants to assess the viability of their ongoing participation. During our experiment, for the case of P11, mentions of repetitive suicide and self-harm were detected. Consequently, an expert contacted the participant, the experiment was suspended for three days, and after a re-evaluation in an outpatient clinic, we resumed the program use with P11.
@@ 6.2. Dialogue Patterns
Participants and MindfulDiary exchanged a total of 4,410 messages (i.e., 2,205 pairs of the AI and participant’s messages) during the field study. Each session consisted of 10.82 messages ($SD=2.70$). The majority of exchanges between the AI and participants were carried on for an exploration of patients’ daily lives and emotions, as well as for casual conversations. In terms of the stage of the conversation, 62% (2,732 messages) of the messages were from Exploration, 30% (1220 messages) for Rapport building, and 6% (282 messages) for Sensitive topic. Only a small amount of messages were accounted for Wrapping up (62 messages) or Not-selected (14 messages).
@@
Category Description Example
Emotional Exploration
Messages that further inquire about the user’s emotional state, mood, and condition.
‘‘How did you feel after meeting her?’’, ‘‘Can you tell me how this situation makes you feel?’’
Activity Exploration
Messages exploring the user’s experiences or activities. These questions delve into more objective and factual content.
‘‘What kind of exercise do you do?’’, ‘‘Lunch with a friend sounds nice. What did you eat?’’
In-depth follow-up
Messages that help delve into the root causes or reasons related to mood, feeling, or specific situations.
‘‘Since when have you felt this overshadowed mood?’’, ‘‘If you feel confused, what might be the reason?’’
Future Plan Exploration
Messages inquiring about how the user reacted to or plans to respond to the events or emotions they mentioned
‘‘That situation must have been challenging. Have you considered any solutions?’’, ‘‘What have you done to alleviate the loneliness, even just a bit?’’
@@
Participants interacting with MindfulDiary conveyed a range of topics (see Table 2). They described a spectrum of emotional states, from negative feelings like exhaustion and anxiety to positive sentiments of pride and joy. Events and activities were recounted, offering insights into their daily routines, such as walking during school times or decreased activity post-vacation. They also shared thoughts and beliefs, sometimes related to current events, revealing patterns linked to mental health, like feelings of exclusion and loneliness. Regarding perceived health status, comments spanned from immediate ailments, such as headaches, to long-term health challenges. Distorted perceptions about their body included content on excessive dieting. Specifically, participants frequently discussed medications, revealing not just their physical reactions but also their perceptions and behaviors toward them. Some expressed concerns over the taste, while others mentioned adverse reactions from intake, like discomfort after swallowing multiple pills at once. Lastly, the realm of relationships & interactions had participants highlighting both the challenges and supports in their interpersonal connections, revealing the significant impact these have on mental well-being, from conflicts and trust issues to moments of affirmation and encouragement.
Category Description Example
Emotional states
The emotions that participants documented in their daily lives encompassed a broad spectrum, ranging from negative and depressive sentiments
“I’m so exhausted, I feel like I’m reaching my limit soon,” “I’m very worried, scared, and anxious”
Events and activities
The mentions of events, tasks, or activities they participated in or witnessed, such as exam periods or travel.
“When I attended school, I got some walking in, but after vacation, I don’t have much reason to go out, leading to a decreased activity level.”
Thoughts and beliefs
The thoughts, values, beliefs, and convictions they usually held. Including characteristic thought patterns related to mental health, such as distorted thinking
“I feel like someone is talking behind my back somewhere; they don’t like my actions and seem to exclude me”
Perceived Health Status
The physical state and health concerns, misconceptions about their health, reactions to medications, revealing their perceptions, behaviors related to drug intake
“I’ll starve and exercise to lose weight!”, “I just took my medicine, but it seems to be getting tasteless.”
Relationships & interactions
The contents about relationships with others. Difficulties stemming from interpersonal relationships, as well as the support and affirmation received from those around them
“I hated seeing my brother being happy. Forcing a cheerful tone also irked me,”
@@ 6.3.1. Broad Conversational Range: The Versatility in Documenting Diverse Interests
Our participants appreciated the LLM’s flexibility and naturalness in responding to various utterances, topics, and situations. Such broad conversational capabilities of the LLM provided participants with a space where they could document a variety of subjects tailored to individual interests and preferences. In our study, participants interacted with the LLM on diverse topics ranging from games, webcomics, novels, and movies to hobbies like making bead bracelets, allowing them to create richer and more personal records. P3 remarked, “AI systems that I have used in the past could only respond to specific words, but it is amazing how this one can respond to all sorts of things.”
@@ 6.3.2. Expanding views: Enriching Entries with Varied Perspectives
Participants also valued the diverse and new perspectives that LLM-generated responses offered, as those helped participants reflect on their daily events and emotions from various angles. This influence helped participants diverge from ruminating on depressive feelings. P12 mentioned, “ Sometimes when you note down emotions, that’s the only thought that comes to mind. Beyond that, I don’t remember much. Since MindfulDiary uses AI, my thoughts flow more easily, and I like it when it asks about different perspectives or topics.”.
@@ 6.3.3. Probing for Depth: Prompt Questions in Detailed Reflection
MindfulDiary’s question-driven journaling process was also valued by participants as it assisted them with the process of daily reflections and documentation. Compared to their past experiences of journaling, where they had to reflect on their daily life by themselves, participants appreciated that MindfulDiary made the journaling process less daunting. P27 said, “Because I have to rely solely on my thoughts when I write alone, I sometimes get stuck. But when I was unsure about how to write, the AI helped me. I liked that part.”. The questions posed by MindfulDiary also guided participants in documenting their daily lives in a more detailed manner by asking their thoughts and feelings about a particular event. Such probing allowed for richer, more in-depth entries. P13 shared, “I used to write diaries on my own and just wrote whatever came to mind. MindfulDiary, however, helped me write in more detail because of the specific questions.”
@@ 6.3.4. Building Narratives: Structuring Daily Reflections with MindfulDiary
MindfulDiary’s capabilities, such as generating contextualized follow-up questions and summarizing the conversation, made the process easier for participants who struggled to organize daily thoughts and events underpsychotherapy (Donnelly and Murray, 1991). In their past experiences, our participants expressed difficulties in journaling because of disjointed thoughts, a lack of clarity in ideas, or inconsistencies in their stories. However, with the support of the LLM in the Mindful Diary, these challenges were addressed, motivating them to record their daily lives persistently. P13 remarked, ” I often had trouble putting sentences In the past, I would worry about writing the next part. But with this tool, I just tell the story of my day, and it seamlessly continues and wraps it up, presenting a well-structured diary entry. That’s its biggest advantage.”

### v2

@@ 5.2.1. MindfulDiary App
Mid-study and Debriefing Interviews. We conducted two 15-minute interviews, after the second and fourth weeks, with each participant to understand their experiences and learn how they used MindfulDiary on a daily basis. Considering the characteristics of depression patients, who may struggle to focus for long periods of time, the interview session was divided into two shorter sessions.
5.2.2. Clinician Dashboard
Most patient participants had a clinical visit during Week 2 through Week 4 of the deployment period. We deployed MindfulDiary’s clinician dashboard to five psychiatrists who are in charge of the participants.
Deployment of Clinician Dashboard in Clinic. We provided instructions to clinicians covering the main components of the clinician dashboard and how to interact with them. To explore the opportunities and limitations of the dashboard, we did not offer explicit instructions for utilizing the clinician dashboard in their workflow. However, we advised psychiatrists to be cautious with the LLM-driven analysis due to potential inaccuracies, emphasizing the importance of verifying data through the interaction logs. The psychiatrists autonomously utilized the clinician dashboard, making sure it didn’t disrupt their current treatment methods and preparation routines.
@@ 5.2.2. Clinician Dashboard
Debriefing Interviews. We interviewed psychiatrists who treated the patient participants to understand how they used the clinician dashboard in clinical settings. We further gathered feedback from the psychiatrists on the opportunities and limitations of MindfulDiary, as well as suggestions for improvements. The interviews with the psychiatrists were conducted offline for about one hour after the deployment study concluded.
Figure 5. Procedure of the four-week field deployment study: A four-week exploration into the utilization of MindfulDiary by outpatient patients, encompassing daily use, and its integration into clinical decision-making. We note that some participants did not have a follow-up visit during the experimental period. The surveys are outside the scope of this work’s investigation.
Timeline diagram illustrating the progression of the field study in the research. Moving from left to right in chronological order: (1) ‘Participant Recruitment,’ (2) ‘introductory session’ (3) ‘4-week MindfulDiary Daily Use’ (4) ‘Surveys & Interviews: Three distinct markers or symbols at the beginning, middle (after two weeks), and end (after four weeks) of the usage period, highlighting the moments when surveys and interviews were conducted. (5) ‘researcher monitoring’
Figure 5. Procedure of the four-week field deployment study: A four-week exploration into the utilization of MindfulDiary by outpatient patients, encompassing daily use, and its integration into clinical decision-making. We note that some participants did not have a follow-up visit during the experimental period. The surveys are outside the scope of this work’s investigation.
@@ 5.3. Analysis
To explore participants’ usage patterns with MindfulDiary, we first conducted a descriptive statistics analysis. To determine any shifts in participants’ adherence over time, we examined weekly writing frequencies using a one-way repeated measures ANOVA (RM-ANOVA) with Greenhouse-Geisser correction. To gain a deeper qualitative understanding of the messages produced by MindfulDiary and interviews with patients and psychiatrists, we used open coding paired with thematic analysis (Braun and Clarke, 2006). For a more in-depth qualitative analysis of the messages produced by MindfulDiary and the interviews with patients and psychiatrists, we employed open coding paired with thematic analysis (Braun and Clarke, 2006). All interviews were audio-recorded and transcribed for this purpose.
The qualitative analysis was conducted by the first author, a PhD student in HCI, who open-coded the interview transcripts and interaction log data through multiple rounds of iteration. Another author who holds a PhD degree in HCI also contributed to this coding process. Following the initial coding, two psychiatrists reviewed the coded data to provide clinical insights and ensure the accuracy of interpretations. Through discussions among the research team, including these diverse perspectives, overarching themes were identified, enhancing the depth and validity of our qualitative findings.
@@ 5.4. Ethical Considerations
Conducting this study, we are fully aware of the inherent risks associated with our research, particularly given the characteristics of participants diagnosed with MDD. To mitigate the risks, we first carefully screened participants, relying on evaluations conducted by psychiatrists. Individuals displaying heightened impulsive tendencies or harboring specific intentions towards self-harm or suicide were excluded from the study. In addition, participants were asked to take the PHQ-9 before interacting with MindfulDiary, along with an additional set of questions probing their recent attempts at self-harm or suicide. If a participant’s response to question number 9 of the PHQ-9, regarding suicidal/self-harm thoughts, scored ‘moderate or higher’ or if any recent suicide attempt was verified, the system pivoted to provide content geared towards alleviating anxiety and reducing stress rather than proceeding with the standard system. In such a case, a real-time alert was also sent to psychiatrists. Lastly, if sensitive themes frequently surfaced in a participant’s input during the study, their interactions with the system were temporarily halted. Psychiatrists subsequently re-evaluated such participants to assess the viability of their ongoing participation. During our experiment, for the case of P11, mentions of repetitive suicide and self-harm were detected. Consequently, an expert contacted the participant, the experiment was suspended for three days, and after a re-evaluation in an outpatient clinic, we resumed the system use with P11.
@@
Figure 6. Overview of daily engagement of participants with MindfulDiary. The colored squares denote the days that participants conversed with MindfulDiary, with darker colors indicating weekend days. The bar charts on the right visualize the total number of days with interaction against the four-week study period. Participants are sorted by the number of days with interaction.
The figure presents an overview of participant engagement with MindfulDiary over a four-week study period. It features a matrix with participants listed on the vertical axis and the days of the study on the horizontal axis. Colored squares indicate days participants interacted with MindfulDiary, with darker shades representing weekends. To the right, bar charts show the total number of days each participant engaged, sorted by their level of interaction. This visualization aims to highlight patterns of usage and identify how engagement varies among participants throughout the study.
Figure 6. Overview of daily engagement of participants with MindfulDiary. The colored squares denote the days that participants conversed with MindfulDiary, with darker colors indicating weekend days. The bar charts on the right visualize the total number of days with interaction against the four-week study period. Participants are sorted by the number of days with interaction.
@@ 6.2. Dialogue Patterns
Participants and MindfulDiary exchanged a total of 4,410 messages (i.e., 2,205 pairs of the AI and participant’s messages) during the field study. Each session consisted of 10.82 messages ($SD=2.70$). Most exchanges between the AI and participants were carried on for an exploration of patients’ daily lives and emotions, as well as for casual conversations. In terms of the stage of the conversation, 62% (2,732 messages) of the messages were from Exploration, 30% (1220 messages) for Rapport building, and 6% (282 messages) for Sensitive topic. Only a small amount of messages were accounted for Wrapping up (62 messages) or not selected (14 messages).
[…]
Table presents a categorization of LLM’s strategies to produce prompts aimed at fostering user reflection. The table is structured into three columns: ’Category’, ’Description’, and ’Example’. ’Emotional Exploration’: This category focuses on probing further into the user’s emotional state, mood, and condition. Sample questions include: “How did you feel after meeting her?” and “Can you tell me how this situation makes you feel?”. ’Activity Exploration’: This strategy explores the user’s experiences or undertakings, leaning towards more factual content. For instance, questions such as “What kind of exercise do you do?” and “Lunch with a friend sounds nice. What did you eat?” fall under this category. ’In-depth follow-up’: This category dives deeper into the root causes or reasons behind the user’s moods, emotions, or particular situations. Example prompts include: “Since when have you felt this overshadowed mood?” and “If you feel confused, what might be the reason?”. ’Future Plan Exploration’: Here, the emphasis is on how the user reacted or plans to tackle the events or emotions they’ve described. Example questions include: “That situation must have been challenging. Have you considered any solutions?” and “What have you done to alleviate the loneliness, even just a bit? Category Description Example Emotional exploration Messages that further inquire about the user’s emotional state, mood, and condition. ‘‘How did you feel after meeting her?’’
‘‘Can you tell me how this situation makes you feel?’’
Activity exploration Messages exploring the user’s experiences or activities. These questions delve into more objective and factual content. ‘‘What kind of exercise do you do?’’
‘‘Lunch with a friend sounds nice. What did you eat?’’
In-depth follow-up Messages that help delve into the root causes or reasons related to mood, feeling, or specific situations. ‘‘Since when have you felt this overshadowed mood?’’
‘‘If you feel confused, what might be the reason?’’
Future plan exploration Messages inquiring about how the user reacted to or plans to respond to the events or emotions they mentioned. ‘‘That situation must have been challenging. Have you considered any solutions?’’
‘‘What have you done to alleviate the loneliness, even just a bit?’’
@@
Participants interacting with MindfulDiary conveyed a range of topics (see Table 4). They described a spectrum of emotional states, from negative feelings like exhaustion and anxiety to positive sentiments of pride and joy. Events and activities were recounted, offering insights into their daily routines, such as walking during school times or decreased activity post-vacation. They also shared thoughts and beliefs, sometimes related to current events, revealing patterns linked to mental health, like feelings of exclusion and loneliness. Regarding perceived health status, comments spanned from immediate ailments, such as headaches, to long-term health challenges. Distorted perceptions about their body included content on excessive dieting. Specifically, participants frequently discussed medications, revealing not just their physical reactions but also their perceptions and behaviors toward them. Some expressed concerns over the taste, while others mentioned adverse reactions from intake, like discomfort after swallowing multiple pills at once. Lastly, the realm of relationships & interactions had participants highlighting both the challenges and supports in their interpersonal connections, revealing their significant impact on mental well-being, from conflicts and trust issues to moments of affirmation and encouragement.
@@ 6.3. Patients’ Perspectives on MindfulDiary
[…]
Table offers a summary of the diverse categories of input messages provided by participants. The table is segmented into three columns: ’Category’, ’Description’, and ’Example’. ’Emotional States’: This category captures the variety of emotions the participants chronicled in their daily interactions. These emotions span a wide array, from negative to depressive feelings. Sample inputs include: “I’m so exhausted, I feel like I’m reaching my limit soon,” and “I’m very worried, scared, and anxious.” ’Events and Activities’: This section includes mentions of events, tasks, or activities participants engaged in or observed, such as academic exams or travel. An example message is: “When I attended school, I got some walking in, but after vacation, I don’t have much reason to go out, leading to a decreased activity level.” ’Thoughts and Beliefs’: Here, participants’ thoughts, values, beliefs, and strong convictions are presented. This category especially encompasses thought patterns linked to mental health like distorted thinking. A representative message is: “I feel like someone is talking behind my back somewhere; they don’t like my actions and seem to exclude me.” ’Perceived Health Status’: This category sheds light on participants’ physical conditions, health misconceptions, reactions to medications, and their perceptions and behaviors concerning medication intake. Example inputs include: “I’ll starve and exercise to lose weight!” and “I just took my medicine, but it seems to be getting tasteless.” ’Relationships & Interactions’: This section brings forward content concerning participants’ relationships with others. It covers challenges rooted in interpersonal dynamics and the support and affirmation they receive from acquaintances. An exemplary message is: “I hated seeing my brother being happy. Forcing a cheerful tone also irked me Category Description Example Emotional states The emotions that participants documented in their daily lives encompassed a broad spectrum, ranging from negative and depressive sentiments. “I’m so exhausted, I feel like I’m reaching my limit soon.”
“I’m very worried, scared, and anxious.”
Events and activities The mentions of events, tasks, or activities they participated in or witnessed, such as exam periods or travel. “When I attended school, I got some walking in, but after vacation, I don’t have much reason to go out, leading to a decreased activity level.” Thoughts and beliefs The thoughts, values, beliefs, and convictions they usually held. Including characteristic thought patterns related to mental health, such as distorted thinking. “I feel like someone is talking behind my back somewhere; they don’t like my actions and seem to exclude me.” Perceived health status The physical state and health concerns, misconceptions about their health, reactions to medications, revealing their perceptions, behaviors related to drug intake. “I’ll starve and exercise to lose weight!”
“I just took my medicine, but it seems to be getting tasteless.”
Relationships & interactions The contents about relationships with others. Difficulties stemming from interpersonal relationships, as well as the support and affirmation received from those around them. “I hated seeing my brother being happy. Forcing a cheerful tone also irked me.”
@@ 6.3.1. Broad Conversational Range: The Versatility in Documenting Diverse Interests
Our participants appreciated the LLM’s flexibility and naturalness in responding to various utterances, topics, and situations. Such broad conversational capabilities of the LLM provided participants with a space where they could document a variety of subjects tailored to individual interests and preferences. In our study, participants interacted with the LLM on diverse topics ranging from games, webcomics, novels, and movies (see 1) to hobbies like Pilates (see 2), allowing them to create richer and more personal records. P3 remarked, “AI systems that I have used in the past could only respond to specific words, but it is amazing how this one can respond to all sorts of things.”
P15 In the afternoon, I watched movies with my family. We watched two films, Asura and The Witness, which were both quite interesting and thrilling.
AI Watching movies with family probably made it more enjoyable. ’Asura’ and ’The Witness’, must have been exciting to watch, since they are both known for their strong stories!
List of dialogues 1 Dialogue snippet from P15
P5 I’ve been dancing to exercise, but I don’t think it’s enough.
AI Dance is a fun and nice form of exercise. But if you feel it’s lacking, Why do you think you feel that way?
P5 I want to exercise by learning from others, like in Pilates.
AI You’ve become interested in Pilates. Can you tell me what makes Pilates interesting to you?
P5 I would like to do Pilates, but I can’t because there’s nowhere to do it.
AI That must be a difficult situation. What about exercising at home? There are many Pilates videos available.
P5 I hadn’t thought of following along with videos.
List of dialogues 2 Dialogue snippet from P5
@@ 6.3.2. Expanding Views: Enriching Entries with Varied Perspectives
Participants also valued the diverse and new perspectives that LLM-generated responses offered, as those helped participants reflect on their struggles, daily events, and emotions from various angles. 3 shows how the system helps the participant to view the challenges of studying from the perspective of the satisfaction felt in gaming. This influence helped participants diverge from ruminating on depressive feelings. P12 mentioned, “Sometimes when you note down emotions, that’s the only thought that comes to mind. Beyond that, I don’t remember much. Since MindfulDiary uses AI, my thoughts flow more easily, and I like it when it asks about different perspectives or topics.”.
AI Such ordinary days accumulate to form our daily lives. What emotions do you feelwhen you study or play games?
P2 I don’t like studying, but I’m working hard dreaming about the college I want to go to. Games are, of course, fun and I feel proud when my character grows.
AI So, studying can be seen as an investment for the future.
P2 Right, studying is an investment for the future.
AI How does it feel when your game character grows? Can you bring that feeling to studying too?
P2 The feeling of my game character growing makes me proud, as if I’ve achieved something. I feel the same pride when I complete my daily study goals.
AI How about continuing to study with that sense of pride? Could it bring the same sense of achievement?
P2 Based on this pride, I plan to focus more on my studies. It won’t be easy, but setting clear goals can lead to faster achievements.
List of dialogues 3 Dialogue snippet from P2
@@ 6.3.3. Probing for Depth: Prompt Questions in Detailed Reflection
MindfulDiary’s question-driven journaling process was also valued by participants as it assisted them with the process of daily reflections and documentation. Compared to their past experiences of journaling, where they had to reflect on their daily life by themselves, participants appreciated that MindfulDiary made the journaling process less daunting. P27 said, “Because I have to rely solely on my thoughts when I write alone, I sometimes get stuck. But when I was unsure about how to write, the AI helped me. I liked that part.”. The questions posed by MindfulDiary also guided participants in documenting their daily lives in a more detailed manner by asking their thoughts and feelings about a particular event (See 4). Such probing allowed for richer, more in-depth entries. P13 shared, “I used to write diaries on my own and just wrote whatever came to mind. MindfulDiary, however, helped me write in more detail because of the specific questions.”
P13 I felt so pathetic and wanted to know why I was acting this way.
AI It must have been really tough to go through such feelings. Does this happen often?
P13 It seems to have gotten worse recently.
AI I see. Do you have any thoughts on why it’s getting worse?
P13 Lately, I’ve been mentally exhausted and stressed a lot.
AI That sounds really difficult. Can you tell me the sources of your stress?
P13 Issues with my dorm, relationships with my friends, and worries about the future.
AI It must be hard with all these concerns overlapping. What’s the most pressing issue for you right now?
P13 These days, I am very concerned about my relationships with friends.
AI I understand your worries about friendships. What is your biggest concern?e
P13 I feel like people are talking behind my back, disliking my actions, and excluding me from gatherings, which makes me feel very lonely.
List of dialogues 4 Dialogue snippet from P13
@@ 6.3.4. Building Narratives: Structuring Daily Reflections with MindfulDiary
MindfulDiary’s capabilities, such as generating contextualized follow-up questions and summarizing the conversation, made the process easier for participants who struggled to organize daily thoughts and events underpsychotherapy (Donnelly and Murray, 1991). In their past experiences, our participants expressed difficulties in journaling because of disjointed thoughts, a lack of clarity in ideas, or inconsistencies in their stories. However, with the support of the LLM in the MindfulDiary, these challenges were addressed, motivating them to record their daily lives persistently. P3 remarked, “ I often had trouble putting sentences In the past, I would worry about writing the next part. But with this tool, I just tell the story of my day, and it seamlessly continues and wraps it up, presenting a well-structured diary entry. That’s its biggest advantage.” (See Section 6.3.4)
P3 The most memorable thing today was my dad eating the cake and food I made yesterday and saying it was delicious.
AI How was the experience of making the cake? Did you enjoy it?
P3 The tough day turned to be enjoyable.
AI Can you describe more about that feeling of happiness?
P3 I felt a bit disappointed and tired while doing it, but when my dad liked it, I felt like I did something great and it made me happy.
[Generated Diary]
‘‘Today was a day without any special worries. The most memorable thing was my dad eating and complimenting the cake and food I made yesterday. That moment turned a somewhat exhausting and difficult day into a joyful one. My father’s praise gave me a great sense of joy and pride. It might be a wrong thought, but I felt really great, like I am an amazing person, which was so nice.’’
List of dialogues 5 Dialogue snippet from P3